# Practical trainable temporal post-processor for multi-state quantum measurement

### v2

@@ E.3 TPP learning of correlated classical noise
In this appendix section, we use a further example to demonstrate the ability of TPP-based learning to extract correlations from measured data, to supplement simulations in Sec. V. Like Sec. V.1, we again consider simulated datasets of measured heterodyne records from a measurement chain of a qubit-cavity-amplifier setup, as in Sec. D.5. Now, however, we consider the excess classical noise added by the measurement process to also possess a component with a colored spectrum (suppressing quadrature labels for clarity):
@@
We immediately see that if the excess classical noise is purely white, the FGDA and TPP exhibit very similar performance: both lie along the dashed line of equal infidelities. However, the situation is very different if the added noise is colored, namely $(\sigma^{\rm P}\!/\sigma^{\rm W})^{2}\neq 0$, and hence has a non-zero correlation timescale. We immediately note that even when the colored noise powers is only a fraction of the white noise power, the TPP-learned filters provide a non-negligible improvement over the standard FGDA scheme using matched filters.
@@ Appendix F Time-shuffled data
As discussed in the previous section, the trained weights $\mathbf{W}$ take the form of Eq. (53),

### v3

@@ E.3 TPP learning of correlated classical noise
In this appendix section, we use a further example to demonstrate the ability of TPP-based learning to extract correlations from measured data, to supplement simulations in Sec. V. Like Sec. V.1, we again consider simulated datasets of measured heterodyne records from a measurement chain of a qubit-cavity-amplifier setup, as in Appendix D.6. Now, however, we consider the excess classical noise added by the measurement process to also possess a component with a colored spectrum (suppressing quadrature labels for clarity):
@@
We immediately see that if the excess classical noise is purely white, the FGDA and TPP exhibit very similar performance: both lie along the dashed line of equal infidelities. However, the situation is very different if the added noise is colored, namely $(\sigma^{\rm P}\!/\sigma^{\rm W})^{2}\neq 0$, and hence has a non-zero correlation timescale. We immediately note that even when the colored noise power is only a fraction of the white noise power, the TPP-learned filters provide a non-negligible improvement over the standard FGDA scheme using matched filters.
@@ Appendix F Time-shuffled data
As discussed in Appendix C, the trained weights $\mathbf{W}$ take the form of Eq. (53),